# Robotic Sorting Systems: Robot Management and Layout Design Optimization

### v3

@@ B.1 Proof of Proposition 1
B.2 Derivation of Maximum Travel Distance in Constraint 5
@@
$v_{1}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\tau_{e}+\frac{v_{VP}}{2c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad\tau_{e}+\frac{v_{VP}}{2c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
@@
$v_{2}(t)=\left\{\begin{aligned} &c_{max}t,\quad t\leq\frac{v_{max}}{c_{max}}\\ &v_{max},\quad\frac{v_{max}}{c_{max}}<t\leq 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}\\ &v_{VP}+c_{max}(2\tau_{e}-t),\quad 2\tau_{e}-\frac{v_{max}-v_{VP}}{c_{max}}<t\leq 2\tau_{e}\\ \end{aligned}\right.$
@@
$d(\tau_{e},v_{VP},v_{max},c_{max})=\left\{\begin{aligned} &d_{1},\quad 2v_{max}-v_{VP}\geq 2\tau_{e}c_{max}\\ &d_{2},\quad 2v_{max}-v_{VP}<2\tau_{e}c_{max}\\ \end{aligned}\right.$
Similarly, the analysis for the acceleration process before a turn is consistent with that of the deceleration process, yielding the same result.
@@
B.3 Derivation of Average Travel Distance in Section 5
@@
Area 3: We first calculate the basic average path length. In accordance with workload balance, the starting and ending points of the path should be equally likely to fall on each workstation. Then add the detour lengths for the two different scenarios multiplied by their respective probabilities, according to the path allocation rules.
@@ B.5 Proof of Proposition 2
The classic penalty successive linear programming (PSLP) enjoys good robustness and convergence properties for large-scale problems (Bazaraa et al.,, 2013). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size. In each iteration $k$, a direction-finding linear program is formulated based on first-order Taylor series approximations to the objective and constraint functions, in addition to appropriate trust region restrictions on the direction components. The subproblem in iteration $k$ are formed as follows:
[…]
@@
Algorithm C-1 The PLSP algorithm to solve model LDP
0:    An initial feasible solution $\omega_{0}$onfidence intervals $0\leq\Delta_{LB}\leq\Delta_{1}$arameters $0\leq\rho_{0}\leq\rho_{1}\leq\rho_{2}\leq 1$, $\phi\leq 1$Large enough constant $\mu$Maximum iterations $N_{m}$

### v4

@@ B.1 Proof of Proposition 1
Assume that the number of vertical and horizontal aisles in the network is $(n_{v},n_{h})$. We illustrate how each VP–cycle reservation in a feasible path can be translated into a set of node–cycle pairs, using the VPs moving along a horizontal aisle as an example.
For a horizontal aisle $a$, we denote the sequence of nodes along the aisle (from entrance node to exit node) as $\mathcal{V}_{aisle}^{(a)}=\{\nu_{1}^{(a)},\nu_{2}^{(a)},\nu_{3}^{(a)},\ldots\}$. The total number of nodes in this aisle is:
$\lvert\mathcal{V}_{ailse}^{(a)}\rvert=2n_{v}+1$
Under the RC-S control scheme, each cycle is divided into four phases, and each phase has a duration of $\tau_{e}$. A VP advances to the next node along its path after each phase. As illustrated in Figure 7 (Section 4), a newly generated VP enters the entrance node of its aisle during the second phase of the cycle in which it is released. In the cycle when it exits the network, it reaches the exit node during the second phase.
Thus, for a VP $p_{0}$ generated in cycle $l_{0}$ along aisle $a$, the set of nodes it occupies in each cycle $l_{i}\in[l_{0},l_{0}+\frac{n_{v}}{2}-1]$ is given by:
$(p_{0},l_{i})\iff\begin{cases}\{(\nu_{1}^{(a)},l_{i}),(\nu_{2}^{(a)},l_{i}),(\nu_{3}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}\\ \{(\nu_{4(l_{i}-l_{0})}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+1}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+2}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+3}^{(a)},l_{i})\},&\text{if }\ l_{0}<l_{i}<l_{0}+\frac{n_{v}}{2}\\ \{(\nu_{2n_{v}}^{(a)},l_{i}),(\nu_{2n_{v}+1}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}+\frac{n_{v}}{2}\\ \emptyset,&\text{otherwise.}\end{cases}$
Similarly, for a VP $p_{0}$ moving along a vertical aisle $a$, we have:
$(p_{0},l_{i})\iff\begin{cases}\{(\nu_{1}^{(a)},l_{i}),(\nu_{2}^{(a)},l_{i}),(\nu_{3}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}\\ \{(\nu_{4(l_{i}-l_{0})}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+1}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+2}^{(a)},l_{i}),(\nu_{4(l_{i}-l_{0})+3}^{(a)},l_{i})\},&\text{if }\ l_{0}<l_{i}<l_{0}+\frac{n_{h}}{2}\\ \{(\nu_{2n_{h}}^{(a)},l_{i}),(\nu_{2n_{h}+1}^{(a)},l_{i})\},&\text{if }\ l_{i}=l_{0}+\frac{n_{h}}{2}\\ \emptyset,&\text{otherwise.}\end{cases}$
$\square$
B.2 Proof of Proposition 2
@@
B.3 Derivation of Maximum Travel Distance in Constraint (7)
@@
[…]
@@
[…]
@@
[…]
Similarly, the analysis for the deceleration process before a turn is consistent with that of the acceleration process, yielding the same result.
@@
B.4 Derivation of Average Travel Distance in Section 5
@@
Area 3: We first calculate the basic average path length. In accordance with workload balance, the starting and ending points of the path should be equally likely to fall on each loading station. Then add the detour lengths for the two different scenarios multiplied by their respective probabilities, according to the path allocation rules.
@@ B.6 Proof of Proposition 3
The classic penalty successive linear programming (PSLP) enjoys good robustness and convergence properties for large-scale problems (bazaraa2013nonlinear). Specifically, PSLP sequentially solves a linearized feasible direction finding subproblem along with the penalty function, and utilizes the concept of trust region (updated at each iteration) to control the step size. In each iteration $k$, a direction-finding linear program is formulated based on first-order Taylor series approximations to the objective and constraint functions, in addition to appropriate trust region restrictions on the direction components. The subproblem in iteration $k$ are formed as follows:
$\textbf{(LP-S)}(\omega_{k},\Delta_{k})$
$\underset{d}{\min}\ \nabla{C_{d}(\omega_{k})}^{T}\cdot d+\mu\cdot(\sum_{i=1}^{2\left|S\right|+2}y_{i})$
@@
To avoid waste of resources, the number of loading station is always equal to the peak number of workers, thus $n_{l}$ is replaced by $\underset{\sigma}{max}\{n_{w}^{\sigma}\}$ in the constraints. The details of solving PLSP is shown in Algorithm C-1. It could be roughly divided into 2 steps in each iteration: (1) Generate LP-S and obtain the step size $d_{k}$, determine whether to stop; (2) Adjust step bounds for the next iteration.
Algorithm C-1 The PLSP algorithm to solve model LDP